you are choosing between two goods. X and Y, and your Marginal utility from each is as shown in the table below. if your income is $9.00 and the price of X and Y are $2 and $1, respectively, what quantities of each will you purchase to maximize utility?
1
Expert's answer
2017-02-06T07:52:10-0500
So, the price of product X = 2, and price of Y is 1 and income is 9, that`s why we have such function as 2*X+Y=9. Then MUX/MUY=PX/PY=2/1, and MUX=2*MUY, that`s why X=3 and Y=3, in this situation such quantities of products are same that maximize utility.